# Infinitely many non-collapsed steady Ricci solitons on complex line bundles

### v1

@@ 5.2. Einstein metrics
We discuss asymptotics for integral curve that represent Ricci-flat metrics in this section. Recall that integral curves that represent Einstein metrics are on the invariant set $\mathcal{E}$, where the function $H-1$ and $Q$ identically vanish.
@@
eventually holds. Since the indexes respectively converge to $-1$ and $-(n-2)$ as $\delta\to 0^{+}$, the inequality
@@
(5.11) $\begin{split}&\left(2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\right)^{\prime}\\ &=\left(2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\right)G+(X_{2}-X_{3})\left(1-4\sqrt{Z_{3}}\right)\\ &\quad+(\sqrt{Z_{2}}-\sqrt{Z_{3}})(-2X_{2}+4(m+1)\sqrt{Z_{3}}-4\sqrt{Z_{2}}+2Z_{1}(\sqrt{Z_{2}}+\sqrt{Z_{3}})).\\ &\geq(X_{2}-X_{3})\left(1-4\sqrt{Z_{3}}\right)\\ &\quad+(\sqrt{Z_{2}}-\sqrt{Z_{3}})(-2X_{2}+4(m+1)\sqrt{Z_{3}}-4\sqrt{Z_{2}}+2Z_{1}(\sqrt{Z_{2}}+\sqrt{Z_{3}})).\\ \end{split}$
@@
If a $\xi(k,\theta,0,0)$ stays in $\mathcal{B}$, the function $X_{1}-X_{2}$ stays positive along the integral curve. Then $Z_{1}$ monotonically increases to $1$ from Proposition 5.5. If $\theta=0$, then the integral curve is in $\mathcal{E}\cap\mathcal{RS}_{\text{FS}}$ and it converges to $p_{1}$. If $\theta>0$, the integral curve converges to $p_{2}$ by Proposition 5.6. ∎
@@
If a $\xi(k,\theta,s_{4},0)$ with $s_{4}>0$ stays in $\mathcal{F}\cap\mathcal{B}$, the first term is eventually positive by Proposition 5.2. By [Chi24, Lemma 4.7], the function $Z_{1}$ converges to 1 and the second term is eventually positive by Proposition 5.1. Since $F_{2}>0$ initially and the function vanishes at $(0,0,0,1,0,0,0,0)$, we reach a contradiction.
@@
Acknowledgements
The author is grateful to McKenzie Wang for valuable discussions on this project. The author also thanks Bohuan Lin and Cheng Yang for helpful discussions on center manifold theory.

### v2

@@ 5.2. Einstein metrics
We discuss asymptotics for integral curve that represent Ricci-flat metrics in this section. Recall that integral curves that represent Einstein metrics are on the invariant set $\mathcal{E}$, where the function H1H-1 and $Q$ identically vanish.
@@
eventually holds. Since the indexes respectively converge to 1-1 and (n2)-(n-2) as $\delta\to 0^{+}$, the inequality
@@
[…]
@@
If a $\xi(k,\theta,0,0)$ stays in $\mathcal{B}$, the function $X_{1}-X_{2}$ stays positive along the integral curve. Then $Z_{1}$ monotonically increases to 11 from Proposition 5.5. If $\theta=0$, then the integral curve is in $\mathcal{E}\cap\mathcal{RS}_{\text{FS}}$ and it converges to $p_{1}$. If $\theta>0$, the integral curve converges to $p_{2}$ by Proposition 5.6. ∎
@@
If a $\xi(k,\theta,s_{4},0)$ with $s_{4}>0$ stays in $\mathcal{F}\cap\mathcal{B}$, the first term is eventually positive by Proposition 5.2. By [Chi26, Lemma 4.7], the function $Z_{1}$ converges to 1 and the second term is eventually positive by Proposition 5.1. Since $F_{2}>0$ initially and the function vanishes at $(0,0,0,1,0,0,0,0)$, we reach a contradiction.